# Geometry is Wavy: Curvature Wave Equations for Generic Affine Connections

### v1

@@ Abstract
Dedicated to Metin Gürses on the occasion of his becoming an emeritus professor after 60 years of research.
@@ I INTRODUCTION
This remarkable equation is purely geometric: it holds in dimensions greater than three without invoking any field equations and follows solely from the differential Bianchi identities. In three dimensions, however, the differential Bianchi identity does not carry independent content beyond the contracted one; consequently, the curvature-wave equation does not arise. The metric signature also plays no essential role at this stage: in Riemannian signature, the same equation becomes a nonlinear elliptic equation rather than a hyperbolic one.
@@
Several important gravitational theories arise as consistent restrictions of the metric-affine framework. General relativity is recovered when both torsion and nonmetricity vanish. Einstein–Cartan theory [8] allows for nonvanishing torsion while preserving metric compatibility, thereby coupling spacetime geometry to the intrinsic spin of matter [9]. Teleparallel theories [10], by contrast, are characterized by vanishing curvature, with gravitation encoded entirely in torsion and/or nonmetricity. In particular, metric teleparallel gravity employs torsion alone [10], symmetric teleparallel gravity employs nonmetricity alone [11], and more general teleparallel models allow for both.
These interrelations are most transparently revealed through the decomposition of a general affine connection into its Levi-Civita [7] part and tensorial corrections associated with torsion and nonmetricity. This geometric landscape shows that gravity may be equivalently described in terms of curvature, torsion, or nonmetricity, depending on the chosen representation. See Figure 1 for a classification of these theories.
@@
Outline of the paper: In Section II we fix conventions and summarize the metric-affine geometric ingredients used throughout (torsion, nonmetricity, and the decomposition of a generic connection into Levi-Civita plus distortion). Section III derives the curvature, the Ricci tensor, and the scalar curvature in terms of these fields and discusses which familiar Riemannian symmetries survive in the generic case. In Section IV we establish the generalized Ricci identities and Bianchi identities needed for consistency checks and for manipulating covariant derivatives. Section V presents the field equations and the main results of the paper, together with the limits that recover standard general relativity. The appendices collect technical derivations and complementary material: Appendix A introduces the decomposition of the generic affine connection into the Levi Civita connection, torsion, and nonmetricity contributions. Appendix B gives the decomposition of the curvature tensors. Appendix C provides the commutator (Ricci) identities; Appendix D derives the generalized Bianchi identities; Appendix E discusses the teleparallel and symmetric-teleparallel specializations [10], and Appendix F summarizes the tetrad and spin–connection [9] formulation and its relation to the metric–affine variables and the last appendix is devoted to a short historical survey.
Figure 1: Schematic classification of gravitational theories according to the geometric properties of spacetime. The table displays the presence or absence of curvature (Riemann tensor $R^{\rho}{}_{\mu\sigma\nu}$), torsion ($T^{\sigma}{}_{\mu\nu}$), and nonmetricity ($Q_{\mu\nu\sigma}$), together with the corresponding form of the affine connection $\Gamma^{\sigma}{}_{\mu\nu}$. Here $^{c}\Gamma^{\sigma}{}_{\mu\nu}$ denotes the Levi–Civita connection, $K^{\sigma}{}_{\mu\nu}$ the contortion tensor associated with torsion, and $L^{\sigma}{}_{\mu\nu}$ the disformation tensor associated with nonmetricity. General relativity is recovered when both torsion and nonmetricity vanish. Einstein–Cartan theory allows for nonvanishing torsion while preserving metric compatibility. Teleparallel theories are characterized by the vanishing of the curvature, with gravitation encoded entirely in torsion and/or nonmetricity.
@@ II CONVENTIONS
In a metric–affine spacetime, the metric $g_{\mu\nu}$ and the affine connection $\Gamma^{\sigma}{}_{\mu\nu}$ are treated as independent geometric objects. The departure from Riemannian geometry is encoded in two tensorial quantities: the nonmetricity
@@
which measures the failure of the connection to preserve the metric under parallel transport, and the torsion
@@
Here $^{c}\Gamma^{\sigma}{}_{\mu\nu}$ denotes the Levi–Civita (Christoffel) connection [7],
@@
$K^{\sigma}{}_{\mu\nu}$ is the contortion tensor constructed from the torsion,
@@
Lowering the first index with the metric, these tensors obey the useful symmetry properties
@@
together with the antisymmetry of the contortion in its first two indices,
@@
Moreover, the contortion tensor is related to the torsion through the identities
@@ III THE CURVATURE TENSORS AND THE BIANCHI IDENTITIES
In terms of the decomposition (10), one can express $R^{\rho}{}_{\mu\sigma\nu}$ as the Levi–Civita [7] curvature plus terms involving $K^{\sigma}{}_{\mu\nu}$ and $L^{\sigma}{}_{\mu\nu}$ (A detailed derivation is given in Appendix B.). We denote the Levi–Civita curvature by $^{c}R^{\rho}{}_{\mu\sigma\nu}$. The Ricci tensor and scalar curvature are defined by the standard contractions
@@
A crucial qualitative difference from the Riemannian case is that, in general, one loses the symmetry under exchange of index pairs,
@@ IV A COMPACT FORM OF THE SECOND BIANCHI IDENTITY AND ITS TRACE
For later convenience we introduce the torsion-dependent and nonmetricity-dependent combinations
@@
Then the second Bianchi identity 24 can be written in the compact form
@@ V WAVE-TYPE EQUATION FOR THE RIEMANN TENSOR
At this point, one may attempt to simplify the last-index divergences $\nabla^{\nu}R_{\gamma\lambda\sigma\nu}$ and $\nabla^{\nu}R_{\gamma\lambda\rho\nu}$ using 28 and (29). In a generic metric-affine geometry, however, these terms do not collapse as neatly as in the Levi-Civita case; this is one of the technical distinctions that makes the metric-affine curvature wave equation structurally richer.
@@
$\quad-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}=0,$ (35)
@@ V.1 Wave equation in Einstein spaces
In this subsection, we specialize the general curvature wave equation (35) to Einstein spaces, i.e. spacetimes whose Ricci tensor is proportional to the metric:
@@
$\quad-\Big(R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}+R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}+R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}\Big)+2\lambda\,R_{\gamma\lambda\sigma\rho}=0.$ (41)
@@ V.2 Wave equation in the torsion- and nonmetricity-free case
In this case, the connection reduces to the Levi-Civita one and, correspondingly, the extra source terms introduced in (26)–(27) disappear:
@@
The Riemann tensor then enjoys its standard Riemannian symmetries, namely
@@
$\quad-\Big(R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}+R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}+R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}+R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}\Big)$
@@
V.3 Wave equation in the nonmetricity-free case ($T^{\mu}{}_{\nu\sigma}\neq 0$)
We next consider the metric-compatible case in which the nonmetricity vanishes,
@@ V.3 Wave equation in the nonmetricity-free case ($T^{\mu}{}_{\nu\sigma}\neq 0$)
Since $Q_{\mu\nu\sigma}=0$, the connection is metric compatible and the covariant derivative preserves the metric, $\nabla_{\mu}g_{\nu\sigma}=0$. Consequently, indices may be raised and lowered freely with $g_{\mu\nu}$ without generating extra nonmetricity terms. In particular, the nonmetricity-dependent tensor introduced in (27) vanishes identically:
@@
$\quad-\Big(R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}+R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}+R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}+R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}\Big)=0.$ (57)
Equation (57) is the metric-compatible torsionful specialization of the general metric-affine curvature wave equation; it will be of particular interest when we further restrict to Einstein-Cartan theory in the next subsection.
V.4 Wave equation in the torsion-free case ($Q^{\mu}{}_{\nu\sigma}\neq 0$)
@@
but, due to the nonmetricity, neither antisymmetry in the first index pair nor the pair-exchange symmetry need hold. In particular,
@@
As a consequence, raising and lowering indices with the metric produces additional terms proportional to $Q_{\mu\nu\sigma}$, and divergences of the Riemann tensor cannot be simplified using the standard Riemannian identities.
@@
$\quad-\Big(R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}+R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}+R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}+R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}\Big)=0.$ (63)
Equation (63) is the torsion-free, nonmetricity including the curvature wave equation. Compared to the metric-compatible case (57), the only additional source term is the divergence of $Y_{\gamma\lambda\nu\rho\sigma}$, which encodes the effect of the nonmetricity on the propagation of the curvature. In the simultaneous limit $T^{\mu}{}_{\nu\sigma}=0$ and $Q_{\mu\nu\sigma}=0$, this equation reduces smoothly to the standard Riemannian curvature wave equation derived in Section V.2.
@@ VI CONCLUSIONS AND FURTHER DISCUSSIONS
In Riemannian geometry [6] equipped with the Levi-Civita connection [7], the curvature of spacetime is not only a measure of gravitational interaction but also a dynamical geometric object. As is well known, the Riemann tensor satisfies a covariant wave equation whose source consists of terms quadratic in curvature [1]. This equation, which follows from the Bianchi identities and the Ricci identity, provides a precise geometric statement of the fact that the curvature propagates: in this sense, the Riemannian geometry is already wavy.
@@
In this work, we have extended this familiar picture to the most general metric-affine setting, where the affine connection possesses the torsion and the nonmetricity, in addition to curvature. Within this framework, we derived a wave-type equation satisfied by the Riemann tensor for a generic connection. The resulting equation exhibits a rich structure: besides the usual curvature–curvature interaction terms known from the Riemannian case, it contains additional contributions driven by the torsion and nonmetricity. These appear both as algebraic couplings to the curvature and as derivative (transport-like) terms, reflecting the fact that torsion and nonmetricity modify the way geometric information propagates across the spacetime.
@@
By analyzing several important limiting cases, we clarified how different geometric structures affect the propagation of curvature. In the purely Riemannian limit, where both the torsion and nonmetricity vanish, our general expression reduces smoothly to the standard curvature wave equation. In Einstein spaces, the equation simplifies further and acquires an effective mass-like term proportional to the cosmological constant. In metric-compatible but torsionful geometries [10], such as Einstein-Cartan theory [8], torsion acts as an additional source and transport mechanism for curvature waves. Conversely, in torsion-free but nonmetric geometries [11], the influence of nonmetricity enters through explicit derivative couplings that have no analogue in Riemannian geometry.
@@
Taken together, these results reinforce a unifying geometric viewpoint: gravity can be described equivalently in terms of the curvature, the torsion, or the nonmetricity [10], but once these structures are present, geometry itself becomes dynamical and wavy. Curvature does not merely encode the static geometric information; it propagates, interacts, and is influenced by the full affine structure of spacetime.
@@
Several directions for further investigation naturally emerge from this work. First, it would be interesting to study the physical interpretation of the additional torsion- and nonmetricity-induced terms in specific models of metric–affine gravity, particularly in relation to spin, dilation, and shear currents of matter. Second, linearizing the general curvature wave equation around the appropriate backgrounds may shed light on the propagation and polarization content of the geometric waves beyond general relativity. Finally, the role of these curvature wave equations in exact solutions, integrability properties, and conservation laws deserves further exploration.
@@
In this appendix, we derive the standard decomposition of a generic affine connection into its Levi–Civita, torsional, and nonmetric contributions. We assume that the spacetime metric $g_{\mu\nu}$ and the affine connection $\Gamma^{\lambda}{}_{\mu\nu}$ are independent.

### v2

@@ Abstract
Dedicated to Metin Gürses on the occasion of his becoming an emeritus professor after 60 years of research and teaching.
@@ I INTRODUCTION
This remarkable equation is purely geometric: it holds in dimensions greater than three without invoking any field equations, and follows solely from the differential Bianchi identities. In three dimensions, however, the differential Bianchi identity does not carry independent content beyond the contracted one; consequently, the curvature-wave equation does not arise. The metric signature also plays no essential role at this stage: in Riemannian signature, the same equation becomes a nonlinear elliptic equation rather than a hyperbolic one.
@@
Several important gravitational theories arise as consistent restrictions of the metric-affine framework. General relativity is recovered when both torsion and nonmetricity vanish. Einstein–Cartan theory [8] allows for nonvanishing torsion while preserving metric compatibility, thereby coupling spacetime geometry to the intrinsic spin of matter [9]. Teleparallel theories [10, 11], by contrast, are characterized by vanishing curvature, with gravitation encoded entirely in torsion and/or nonmetricity. In particular, metric teleparallel gravity employs torsion alone [10], symmetric teleparallel gravity employs nonmetricity alone [12, 13, 14, 15], and more general teleparallel models allow for both.
These interrelations are most transparently revealed through the decomposition of a general affine connection into its Levi-Civita [7] part and tensorial corrections associated with torsion and nonmetricity. This geometric landscape shows that gravity may be equivalently described in terms of curvature, torsion, or nonmetricity, depending on the chosen representation. See Figure 1 for a classification of these theories. [^1] We owe the last line of the table to F.W. Hehl, see the nice exposition [16]
@@
Outline of the paper: In Section II we fix conventions and summarize the metric-affine geometric ingredients used throughout (torsion, nonmetricity, and the decomposition of a generic connection into Levi-Civita plus distortion). Section III derives the curvature, the Ricci tensor, and the scalar curvature in terms of these fields and discusses which familiar Riemannian symmetries survive in the generic case. In Section IV we establish the generalized Ricci identities and Bianchi identities needed for consistency checks and for manipulating covariant derivatives. Section V presents the field equations and the main results of the paper, together with the limits that recover standard general relativity. The appendices collect technical derivations and complementary material: Appendix A introduces the decomposition of the generic affine connection into the Levi Civita connection, torsion, and nonmetricity contributions. Appendix B gives the decomposition of the curvature tensors. Appendix C provides the commutator (Ricci) identities; Appendix D derives the generalized Bianchi identities; Appendix E discusses the teleparallel and symmetric-teleparallel specializations [10], while Appendix F summarizes the tetrad and spin connection [9] formulation and its relation to the metric-affine variables. The last appendix is devoted to a short historical survey.
Figure 1: Schematic classification of gravitational theories according to the geometric properties of spacetime. The table displays the presence or absence of curvature (Riemann tensor $R^{\rho}{}_{\mu\sigma\nu}$), torsion ($T^{\sigma}{}_{\mu\nu}$), and nonmetricity ($Q_{\mu\nu\sigma}$), together with the corresponding form of the affine connection $\Gamma^{\sigma}{}_{\mu\nu}$. Here $^{c}\Gamma^{\sigma}{}_{\mu\nu}$ denotes the Levi–Civita connection, $K^{\sigma}{}_{\mu\nu}$ the contorsion tensor associated with torsion, and $L^{\sigma}{}_{\mu\nu}$ the disformation tensor associated with nonmetricity. General relativity is recovered when both torsion and nonmetricity vanish. Einstein–Cartan theory allows for nonvanishing torsion while preserving metric compatibility. Teleparallel theories are characterized by the vanishing of the curvature, with gravitation encoded entirely in torsion and/or nonmetricity.
@@ II CONVENTIONS
In a metric-affine spacetime, the metric $g_{\mu\nu}$ and the affine connection $\Gamma^{\sigma}{}_{\mu\nu}$ are treated as independent geometric objects. The departure from Riemannian geometry is encoded in two tensorial quantities: the nonmetricity
@@
which measures the failure of the connection to preserve the metric under parallel transport and the torsion
@@
Here $^{c}\Gamma^{\sigma}{}_{\mu\nu}$ denotes the Levi-Civita (Christoffel) connection [7],
@@
$K^{\sigma}{}_{\mu\nu}$ is the contorsion tensor constructed from the torsion,
@@
Lowering the first index with the metric, these tensors obey useful symmetry properties
@@
together with the antisymmetry of the contorsion in its first two indices,
@@
Moreover, the contorsion tensor is related to the torsion through the identities
@@ III THE CURVATURE TENSORS AND THE BIANCHI IDENTITIES
In terms of the decomposition (10), one can express $R^{\rho}{}_{\mu\sigma\nu}$ as the Levi-Civita [7] curvature plus terms involving $K^{\sigma}{}_{\mu\nu}$ and $L^{\sigma}{}_{\mu\nu}$. (A detailed derivation is provided in Appendix B.) We denote the Levi-Civita curvature by $^{c}R^{\rho}{}_{\mu\sigma\nu}$. The Ricci tensor and scalar curvature are defined by the standard contractions
@@
A crucial qualitative difference from the Riemannian case is that, in general, one loses the symmetry under the exchange of index pairs,
@@ IV A COMPACT FORM OF THE SECOND BIANCHI IDENTITY AND ITS TRACE
For later convenience, we introduce the torsion-dependent and nonmetricity-dependent combinations
@@
Then, the second Bianchi identity 24 can be written in a compact form
@@ V WAVE-TYPE EQUATION FOR THE RIEMANN TENSOR
At this point, one may attempt to simplify the last-index divergences $\nabla^{\nu}R_{\gamma\lambda\sigma\nu}$ and $\nabla^{\nu}R_{\gamma\lambda\rho\nu}$ using (28) and (29). In a generic metric-affine geometry, however, these terms do not collapse as neatly as in the Levi-Civita case; this is one of the technical distinctions that makes the metric-affine curvature wave equation structurally richer.
@@
$\Box R_{\gamma\lambda\rho\sigma}+\nabla_{\rho}\nabla^{\nu}R_{\gamma\lambda\sigma\nu}-\nabla_{\sigma}\nabla^{\nu}R_{\gamma\lambda\rho\nu}+\nabla^{\nu}\!\left(X_{\gamma\lambda\nu\rho\sigma}+Y_{\gamma\lambda\nu\rho\sigma}\right)$ $+T^{\nu k}{}_{\rho}\,\nabla_{\nu}R_{\gamma\lambda\sigma k}-T^{\nu k}{}_{\sigma}\,\nabla_{\nu}R_{\gamma\lambda\rho k}$ $+R^{k}{}_{\gamma\rho}{}^{\nu}R_{k\lambda\sigma\nu}+R^{k}{}_{\lambda\rho}{}^{\nu}R_{\gamma k\sigma\nu}+R^{k}{}_{\sigma\rho}{}^{\nu}R_{\gamma\lambda k\nu}+R^{\kappa}\ _{\rho}R_{\gamma\lambda\sigma\kappa}$ $-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}-R^{\kappa}\ _{\sigma}R_{\gamma\lambda\rho\kappa}$ $=0,$ (35)
@@ V.1 Wave equation in Einstein spaces
In this subsection, we specialize the general curvature wave equation (35) to Einstein spaces, i.e., spacetimes whose Ricci tensor is proportional to the metric:
@@
$\quad-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}+2\lambda\,R_{\gamma\lambda\sigma\rho}=0.$ (41)
@@ V.2 Wave equation in the torsion- and nonmetricity-free case
In this case, the connection reduces to the Levi-Civita one, and correspondingly, the extra source terms introduced in (26)–(27) disappear:
@@
The Riemann tensor then enjoys its standard Riemannian symmetries, namely,
@@
$\quad-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}$
@@
V.3 Wave equation in the nonmetricity-free case
We next consider the metric-compatible case in which nonmetricity vanishes,
@@ V.3 Wave equation in the nonmetricity-free case
Since $Q_{\mu\nu\sigma}=0$, the connection is metric compatible, and the covariant derivative preserves the metric, $\nabla_{\mu}g_{\nu\sigma}=0$. Consequently, indices may be raised and lowered freely with $g_{\mu\nu}$ without generating extra nonmetricity terms. In particular, the nonmetricity-dependent tensor introduced in (27) vanishes identically:
@@
$\quad-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}=0.$ (57)
Equation (57) is the metric-compatible, torsionful specialization of the general metric-affine curvature wave equation; it will be of particular interest when we further restrict to the Einstein-Cartan theory in the next subsection.  [^8] For electromagnetic waves propagating on a background with nonvanishing torsion, see [17].
V.4 Wave equation in the torsion-free case
@@
but, due to nonmetricity, neither antisymmetry in the first index pair nor pair-exchange symmetry needs to hold. In particular,
@@
As a consequence, raising and lowering indices with the metric produces additional terms proportional to $Q_{\mu\nu\sigma}$, and the divergences of the Riemann tensor cannot be simplified using the standard Riemannian identities.
@@
$\quad-R^{k}{}_{\gamma\sigma}{}^{\nu}R_{k\lambda\rho\nu}-R^{k}{}_{\lambda\sigma}{}^{\nu}R_{\gamma k\rho\nu}-R^{k}{}_{\rho\sigma}{}^{\nu}R_{\gamma\lambda k\nu}-R^{k}{}_{\ \sigma}\,R_{\gamma\lambda\rho k}=0.$ (63)
Equation (63) is the torsion-free, nonmetricity including the curvature wave equation. Compared to the metric-compatible case (57), the only additional source term is the divergence of $Y_{\gamma\lambda\nu\rho\sigma}$, which encodes the effect of nonmetricity on the propagation of curvature. In the simultaneous limit $T^{\mu}{}_{\nu\sigma}=0$ and $Q_{\mu\nu\sigma}=0$, this equation reduces smoothly to the standard Riemannian curvature wave equation derived in Section V.2. In addition to the given constraints, namely, the vanishing of torsion and the nonvanishing of nonmetricity, one may further assume the vanishing of the Riemann tensor, which leads to Symmetric Teleparallel Gravity (STGR). Moreover, although general relativity has no local degrees of freedom in three dimensions, STGR does possess them [18].
@@ VI CONCLUSIONS AND FURTHER DISCUSSIONS
In Riemannian geometry [6] equipped with the Levi-Civita connection [7], the curvature of spacetime is not only a measure of gravitational interaction but also a dynamical geometric object. As is well known, the Riemann tensor satisfies a covariant wave equation whose source consists of terms quadratic in curvature [1]. This equation, which follows from the Bianchi identities and the Ricci identity, provides a precise geometric statement of the fact that curvature propagates: in this sense, Riemannian geometry is already wavy.
@@
In this work, we have extended this familiar picture to the most general metric-affine setting, where the affine connection possesses torsion and nonmetricity, in addition to curvature. Within this framework, we derived a wave-type equation satisfied by the Riemann tensor for a generic connection. The resulting equation exhibits a rich structure: besides the usual curvature–curvature interaction terms known from the Riemannian case, it contains additional contributions driven by the torsion and nonmetricity. These appear both as algebraic couplings to the curvature and as derivative (transport-like) terms, reflecting the fact that torsion and nonmetricity modify the way geometric information propagates across spacetime.
@@
By analyzing several important limiting cases, we clarified how different geometric structures affect the propagation of curvature. In the purely Riemannian limit, where both the torsion and nonmetricity vanish, our general expression reduces smoothly to the standard curvature wave equation. In Einstein spaces, the equation simplifies further and acquires an effective mass-like term proportional to the cosmological constant. In metric-compatible but torsionful geometries [10], such as Einstein-Cartan theory [8], torsion acts as an additional source and transport mechanism for curvature waves. Conversely, in torsion-free but nonmetric geometries [12], the influence of nonmetricity enters through explicit derivative couplings that have no analog in Riemannian geometry.
@@
Taken together, these results reinforce a unifying geometric viewpoint: Gravity can be described equivalently in terms of curvature, torsion, or nonmetricity [10], but once these structures are present, geometry itself becomes dynamic and wavy. Curvature does not merely encode static geometric information; it propagates, interacts, and is influenced by the full affine structure of spacetime.
@@
Several directions for further investigation naturally emerge from this work. First, it would be interesting to study the physical interpretation of the additional torsion- and nonmetricity-induced terms in specific models of metric-affine gravity, particularly in relation to spin, dilation, and shear currents of matter. One can find works in the literature along this direction [19, 20] Second, linearizing the general curvature wave equation around the appropriate backgrounds may shed light on the propagation and polarization content of the geometric waves beyond general relativity. Finally, the role of these curvature wave equations in exact solutions, integrability properties, and conservation laws deserves further exploration.
@@
In this appendix, we derive the standard decomposition of a generic affine connection into its Levi-Civita, torsional, and nonmetric contributions. We assume that the spacetime metric $g_{\mu\nu}$ and the affine connection $\Gamma^{\lambda}{}_{\mu\nu}$ are independent.